OpenStudy (anonymous):

what is the coordinate in quadrant 1 for 3y^2-5x^2=7 and 3y^2-x^2=23

OpenStudy (anonymous):

I would use elimination to eliminate the y^2 term and solve for the x^2 term and then x.

OpenStudy (anonymous):

can you help me a little more than that.

OpenStudy (anonymous):

\[3y^2-5x^2=7\] \[-\left( 3y^2-x^2=23 \right)\] This will eliminate the y^2 term now solve for x^2 then take the squareroot to solve for x...Then find the value of y by substituting the x value into either original equation.

OpenStudy (anonymous):

so what would be the final answer then?

OpenStudy (anonymous):

Show me your work so far...

OpenStudy (anonymous):

i dont know where to start so i have no work

OpenStudy (anonymous):

So you just wanted me to give you the answer?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

The name of this website is openstudy.com not givemeanswers.com. If you want help, I'd be glad to. But if you just want answers, try cheaters.com.

OpenStudy (anonymous):

ok thank you anyways.

OpenStudy (anonymous):

By the way I think I explained how you could start.